Use the following information to answer the next three questions. Consider the following cash flows: Year Cash Flow O $2,400,What is the payback period? Round your answer to two decimal places.

What is the net present value (NPV) rounded to the nearest dollar?

What is the internal rate of return?

Please show all work. Thank you.
Answer #1

Cost = 2400000

In 3 years this project will generate 1850000.

Remaining amount to recover = 2400000 - 1850000 = 550000

Time required to recover remaining amount = 550000 / 900000 = 0.61

Payback Period = 3 + 0.61 = 3.61 years Answer

NPV = Present value of all future cash inflow - initial outlay

NPV = 780000/(1+0.178)^1 + 510000/(1+0.178)^2 + 560000/(1+0.178)^3 + 900000/(1+0.178)^4 - 2400000

NPV = -$560399.17 Answer

IRR is the rate for which NPV equals zero.

0 = 780000/(1+IRR)^1 + 510000/(1+IRR)^2 + 560000/(1+IRR)^3 + 900000/(1+IRR)^4 - 2400000

We will use heat and trial method to get that value for which above equation satisfy.

IRR = 5.52% Answer

Please let me know in case you have any queries and I will be happy to assist you.